OpenStudy (anonymous):

Simplify the following expression; express in terms of i. (square root sign) -81

OpenStudy (saifoo.khan):

9i

OpenStudy (lalaly):

\[\sqrt{-81} = \sqrt{-1 \times 81} = \sqrt-1 \times \sqrt81 = 9i\]